OpenStudy (anonymous):

4/y+5=3/y-5 please help me solve.

OpenStudy (anonymous):

add 5 to both side so 4/y +10 = 3/y

OpenStudy (anonymous):

4/y-3/y=-10 1/y(4-3)=-10

OpenStudy (anonymous):

1/y=-10 y=-1/10